Assessment activity: Chapter 05- Making Automobile and Housing Decisions buy analysis- Part 2 10. Rent versus Aa Aa Which Is Better: To Rent or To Buy? The decision of whether to rent or buy housing is a personal decision that is based on both your lifestyle and your finances. While most financial experts argue that the financial aspect of the decision is important, it is also important not to base your rent-or-buy decision solely on the numbers. Your personal needs and housing market conditions are impórtant. However, it is still necessary to perform the financial analysis. Megan and her husband are trying to decide whether to rent or to purchase a new house. After looking for several months, they've narrowed their choice down to one particular house, and the builder is willing to lease or sell-depending upon the preference of the buyer. To perform a rent-or-buy analysis, Megan and her husband have collected the following information: If they rent, the builder will require monthly rental payments of $1,200 and a security deposit equal to two months of rent. Since they want to be protected against the possible loss of their possessions, they will purchase a renters' policy of $200 every six months, while a more comprehensive homeowners' policy will cost 0.5% of the home's value per Complete a rent-or-buy analysis worksheet to determine the total cost of renting and the total cost of purchasing Megan and her husband's prospective house and then recommend their best strategy. To complete the worksheet, enter the appropriate values in their corresponding blanks and round each value to the nearest whole dollar.
tivity: Chapter 05- Making Automobile and Housing Decisions RENT-OR-BUY ANALYSIS FOR HOUSING $ Amount COST OF RENTING Security deposit Annual rental cost Renter's insurance Opportunity cost on security deposit Total Annual Cost of Renting: COST OF BUYING Monthly mortgage payment Annual mortgage payments Property taxes Homeowner's insurance Maintenance expenses Opportunity cost of down payment and closing costs Total costs Less Reduction of loan principal Tax savings on mortgage interest deduction Tax saving on property tax deduction Total deductions Annual after-tax cost of homeownership Estimated annual appreciation in home value Total Annual Cost of Purchasing:
year. Money used to fund the house's security deposit could otherwise be invested to earn 5 % per year after taxes. Funds expended for a home's down payment and closing costs also incur an opportunity cost... If the house is purchased, it will cost $200,000 and will require a 20% down payment. The loan willl carry an interest rate of 6 %, a term of 30 years, and monthly payments of $959. The closing costs associated with the house's mortgage will be $3,500.
The property taxes and the maintenance and repair expenses on the house are estimated to be 4% and 2 % of the house's total price, respectively. Your ordinary income is taxed at the rate of 28%, and you'll be willing to itemize your tax deductions in the event that you purchase your new home. Financial publications report that home values are expected to increase by 3 % this year due to inflation.
Answer #1

Rent or Buy analysis for Housing

1. Cost of renting

(All figures are in annual)

Security Deposit = $1200 * 2= $2,400

Annual rental cost = $1200 * 12 = $14,400

Renter's insurance = $200 * 2 = $1,200

Opportunity cost on Security deposit = $2,400 * 5% = $120

Total Annual cost of renting = $2,400 + $14,400 + $1,200 + $120 = $18,120.

Cost of Buying

(All figures are in annual)

Monthly Mortgage payment = $959

Annual Mortgage payment = $959 * 12 = $11,508

Property Taxes = 4% * $200,000 = $8,000

Homeowner's Insurance = 0.5% * $200,000 = $1,000

Maintenance expenses = 2% * $200,000 = $4,000

Opportunity cost of down payment and closing costs = 20%* $200,000* 5% + $3500 * 5% = $2,175

Total cost = $11,508 + $8,000 + $1,000 + $4,000 + $2,175 = $26,683

Less

Reduction of loan principal = $1,964 (approximate as per amortization table for first year)

Tax savings on mortgage interest deduction = ($11,508 - $1,964) * 28%= 9,544* 28% = $2,672

Tax saving on property tax deduction = $8,000 * 28% = $2,240

Total deductions = $2,672 + $2,240 = $4,912

Annual after tax cost of ownership = $26,683 - $1,964 - $4,912 = $19,807

Estimated annual appreciation in home value = 3% * $200,000 = $6,000

Total Annual cost of Purchasing = $19,807 - $6,000 = $13,807.

Considering annual cost of purchasing ($13,807) is lower as compared to annual lost of renting ($18,120) hence purchasing house is a better strategy.

If you liked my answer, please give like/ thumbs up. Thank you!